OpenStudy (anonymous):

simplify √4/9 which i got 2/3

OpenStudy (anonymous):

is the entire thing under the root? if so your correct

OpenStudy (anonymous):

nd give medals

OpenStudy (anonymous):

\[\sqrt{4/9}=2/3\] you are correct

OpenStudy (anonymous):

thanks